OpenStudy (anonymous):

there were 80.8 voters one year then the next there was 117.5. Find the percent change in the number of voters.

OpenStudy (anonymous):

[(117.5 - 80.8)/80.8] x 100

OpenStudy (anonymous):

117.5=80.8*(1+p/100) 1+p/100=1.454 p/100=0.454 p=45.4 %

OpenStudy (anonymous):

You take the change minus the original number, quantity divided by the original number. For %, you then multiply by 100.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

With either method, you get 45.4 %. With my method, you get it in one step.